Mthsc 881 Mathematical Statistics 2.30-3.45 MW, Martin M103

Dr. K.B. Kulasekera Martin O 322; 656-5231 Office Hrs: 1.30-3.00 pm T,TR, 4.00-4.45 MW or by appt. e-mail : kk@math.clemson.edu

Text: Modern Mathematical Statistics by Dudewicz and Mishra and Class Handouts. This course is designed to provide you with a good understanding of the theory and moti-

vation behind some well known and widely used statistical techniques. In particular we discuss Likelihood Principle, Sufficiency and Completeness, Likelihood Ratio Tests and Invariance Principle.

Syllabus

1. Limit Laws (Chapter 6)

2. Point Estimation (Chapters 7& 8)

3. Tests of Hypotheses (Chapter 9)

4. Confidence Intervals (Chapter 10)

5. Location and Scale Families and Invariance (Class Handouts)

Grading:

Home Work...................................................40% In Class tests(2).......................................... 40% Final Exam (In Class)...................................20%

Grades: ≥ 90% -A;[80, 90)% -B;[70 − 80)%-C.

1. All exams are open book. We may have to meet on two Friday afternoons for in class tests

2. All drop/add procedures are the responsibility of individual student.

3. Home work problems should be solved independently as much as possible.

4. If I am more than 15 minutes late, the class is considered cancelled.